OpenStudy (anonymous):

3 + a + b = a + 3 + b is an example of which algebraic property?

OpenStudy (mokeira):

Distributive: 3(a + b) = 3a + 3b Commutative (Addition): a + b + c = c + a + b Associative (Addition): (a + b) + c = a + (b + c) Symmetric (Addition): a = b, b = a

OpenStudy (mokeira):

so i guess it is Commutative (Addition): a + b + c = c + a + b